# Towards Trustworthy Federated Learning with Untrusted Participants

### v1

@@ Abstract
Resilience against malicious parties and data privacy are essential for trustworthy distributed learning, yet achieving both with good utility typically requires the strong assumption of a trusted central server. This paper shows that a significantly weaker assumption suffices: each pair of workers shares a randomness seed unknown to others. In a setting where malicious workers may collude with an untrusted server, we propose CafCor, an algorithm that integrates robust gradient aggregation with correlated noise injection, leveraging shared randomness between workers. We prove that CafCor achieves strong privacy-utility trade-offs, significantly outperforming local differential privacy (DP) methods, which do not make any trust assumption, while approaching central DP utility, where the server is fully trusted. Empirical results on standard benchmarks validate CafCor’s practicality, showing that privacy and robustness can coexist in distributed systems without sacrificing utility or trusting the server.
Differential Privacy, Byzantine, Robust, Distributed Optimization, Federated Learning
@@ 1 Introduction
Differential privacy (DP) (Choudhury et al., 2019; Hu et al., 2020; Noble et al., 2022) and robustness (Blanchard et al., 2017; Yin et al., 2018; Karimireddy et al., 2021) have been extensively studied in isolation in distributed machine learning. However, their intersection, which is crucial for building trustworthy learning systems, remains underexplored. Recent work shows that combining local DP (LDP), which assumes no trust, with robustness incurs a fundamental utility cost (Allouah et al., 2023b). This reinforces earlier findings that ensuring privacy under LDP, even without malicious workers, is inherently prohibitive in terms of utility (Duchi et al., 2013). In contrast, the central DP (CDP) model, where the server is fully trusted, enables significantly better utility than LDP, particularly as the number of workers increases (Liu et al., 2021; Hopkins et al., 2022). These findings highlight the trade-offs between trust, privacy, and robustness in distributed learning, underscoring the need for alternative approaches that mitigate these limitations.
@@ 1.2 Related Work
Most works combining DP and robustness focus on local privacy models, often only achieving per-step privacy guarantees and suboptimal learning performance (Guerraoui et al., 2021; Zhu & Ling, 2022; Xiang & Su, 2022; Ma et al., 2022). Others reveal a three-way trade-off between privacy, robustness, and utility for discrete distribution estimation in non-interactive settings (Cheu et al., 2021; Acharya et al., 2021) and more general interactive learning tasks (Allouah et al., 2023b). In the CDP model where the server is fully trusted, the privacy-utility trade-off is far superior to the LDP setting (Duchi et al., 2013, 2018). Prior work on centralized robust mean estimation with DP (Liu et al., 2021; Hopkins et al., 2022; Alabi et al., 2023) achieves near-optimal error bounds, disentangling the three-way trade-off by matching the optimal sample complexity in both private and robust settings. Empirical studies confirm this advantage in standard federated learning tasks (Xie et al., 2023).
Intermediate threat models between local and central DP, such as those leveraging cryptographic primitives or anonymization via shuffling (Jayaraman et al., 2018; Erlingsson et al., 2019; Cheu et al., 2019), can approach the privacy-utility trade-off of central DP but often incur high computational costs, rely on trusted entities, or lack robustness. Correlated noise schemes have also been explored for privacy without a trusted server; early methods like secret sharing (Shamir, 1979) ensure input security but do not privatize the output, leaving it vulnerable to inference attacks (Melis et al., 2019). More refined approaches (Imtiaz et al., 2021) combine correlated and uncorrelated Gaussian noise to protect averages, while decentralized extensions remain limited to averaging tasks (Sabater et al., 2022) and lack robustness (Allouah et al., 2024). We also note that (Sabater et al., 2022) consider malicious workers but do not provide utility guarantees in their presence, and only use cryptographic checks, e.g., to verify boundedness of all messages. Recently, Gu et al. (2023) considered malicious workers and an untrusted server but relied on computationally intensive tools like zero-knowledge proofs and secure aggregation, whereas our method avoids these complexities. Moreover, Gu et al. (2023) do not quantify the utility cost of their solution.
@@ 2.3 Assumptions
Our results are derived under standard assumptions. Data heterogeneity is modeled via Assumption 2.1 (Allouah et al., 2023b), which subsumes standard assumptions in federated learning (Karimireddy et al., 2020) by bounding gradient variance in all directions. To establish the convergence guarantees of CafCor, we adopt the standard Assumption 2.2 on the variance of stochastic gradients. Assumption 2.3 ensures bounded gradients at any data point, a common requirement in differentially private optimization to avoid complications due to clipping (Agarwal et al., 2018).
@@ 4.1 Privacy Analysis
The proof of Theorem 4.1, deferred to Appendix A, relies upon computing the Rényi divergence induced by the correlated noise scheme, before converting back to DP following standard conversion results (Mironov, 2017). Indeed, we derive the exact expression of the aforementioned Rényi divergence in Lemma A.4, which we use directly in our Rényi-based privacy accountant for SecLDP in practice. This exact computation was not feasible in previous related works, as they either considered different communication topologies (Allouah et al., 2024) or did not analyze privacy through the lens of Rényi divergences (Sabater et al., 2022).
@@ 4.2 Utility Analysis
Proposition 4.1 establishes an important property satisfied by CAF. The above corresponds to the high-dimensional robustness criterion studied by Allouah et al. (2023b), shown to be stronger than existing ones in the distributed learning literature, e.g., that of Karimireddy et al. (2022). Furthermore, we recall from Allouah et al. (2023b) that standard robust aggregations, such as the trimmed mean and median (Yin et al., 2018), can only achieve a weaker variant of Proposition 4.1, where the right-hand side of the inequality involves the trace instead of the maximum eigenvalue operator. Since the trace of a $d\times d$ matrix can be up to $d$ times its largest eigenvalue, standard robust aggregations only satisfy Proposition 4.1 with an additional multiplicative factor of $d$ in $\kappa$ (4.1), leading to a looser robustness bound.
Moreover, the so-called SMEA aggregation of Allouah et al. (2023b) also satisfies Proposition 4.1 up to constants. However, SMEA is computationally impractical as its time complexity grows exponentially with $f$. In contrast, CAF is efficient, running in polynomial time $\mathcal{O}(f(nd^{2}+d^{3}))$. Indeed, at each iteration, at least one input ($i\in\operatorname*{argmax}_{j\in[n]}\tau_{j}$) is assigned a weight of zero. Thus, CAF terminates within at most $2f$ iterations, each requiring the computation of the top eigenvector and the formation of the empirical covariance matrix.To further improve efficiency, we implement a CAF variant with reduced complexity $\mathcal{O}(fnd\log{d})$ by approximating the top eigenvector using the power method. This approach avoids explicit covariance matrix formation and instead relies on matrix-vector products. We adopt this more efficient variant in our experiments in Section 5, as it can be easily shown to satisfy Proposition 4.1 up to constants.
@@
Our result recovers the state-of-the-art convergence analysis in robust optimization in the local privacy model (Allouah et al., 2023b) when setting $\sigma_{\mathrm{cor}}=0$. The key difference with the latter analysis is the role of the correlated noise magnitude $\sigma_{\mathrm{cor}}$, which appears in the dominating term inside $\overline{\sigma}$, in both strongly convex and non-convex settings. Moreover, our convergence rates exhibit the standard dependence on the number of iterations in smooth stochastic optimization: $\mathcal{O}(\tfrac{1}{T})$ for the strongly convex case and $\mathcal{O}(\tfrac{1}{\sqrt{T}})$ for the non-convex case. The presence of a non-vanishing term in $T$ is fundamental in robust optimization in heterogeneous data settings (Karimireddy et al., 2022). Finally, our upper bound in the strongly convex case also applies to non-convex loss functions that satisfy the Polyak-Lojasiewicz inequality (Karimi et al., 2016).
@@ 6 Conclusion
First, we remark that Algorithm 2 terminates after at most $2f$ iterations. Indeed, at each iteration, at least one weight $c_{i},i\in[n]$, becomes zero—this is true for any $i\in[n]$ such that $i\in\operatorname*{argmax}_{j\in[n]}\left\langle\hat{v}_{c},\,x_{j}-\hat{\mu}_{c}\right\rangle^{2}$—for the remaining iterations. Therefore, since the weights $c_{i},i\in[n],$ are all in $[0,1]$, it is guaranteed that $\sum_{i=1}^{n}c_{i}\leq n-2f$ after $2f$ iterations.
@@ C.2 Proof Outline
Our analysis of CafCor (Algorithm 1), is inspired from standard analyses in non-private robust distributed learning (Karimireddy et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023b). We include all proofs for completeness.
@@
The dimension factor $d$ due to correlated and uncorrelated noises is divided by $n-f$, which would not have been possible without leveraging the Gaussian nature of the noise. To do this, we use a concentration argument on the empirical covariance matrix of Gaussian random variables, stated in Lemma C.6. Moreover, the correlated noise magnitude $\sigma_{\mathrm{cor}}^{2}$ has a multiplicative factor $n$, which essentially reflects the fact that a correlated noise term is added per worker, for each worker. Finally, the remaining term $G_{\mathrm{cov}}^{2}$ of the upper bound is only due to data heterogeneity, and also appears in Allouah et al. (2023b). Similar to the aforementioned work, an important distinction from Karimireddy et al. (2022) is that $G_{\mathrm{cov}}^{2}$ is a tighter bound on heterogeneity, compared to $G^{2}$ the bound on the average squared distance from 2.1. This is because the drift $\Delta_{t}$ is not an average squared distance, but rather a bound on average squared distances of every projection on the unit ball. Controlling this quantity requires a covering argument (stated in Lemma C.4).
@@
We prove Theorem 4.2 in the strongly convex case in Section C.3.1, and in the non-convex case in Section C.3.2. These proofs follow along the lines of Allouah et al. (2023b), and we include them here for completeness.
@@ C.6 Proof of Supporting Lemmas
Before proving Lemmas C.1, C.2 and C.3 in Sections C.6.2, C.6.3 and C.6.4, respectively, we first present some additional results in Section C.6.1 below. Most of the proofs follow along the lines of Allouah et al. (2023b), and we include them here for completeness.

### v2

@@ Abstract
Resilience against malicious participants and data privacy are essential for trustworthy federated learning, yet achieving both with good utility typically requires the strong assumption of a trusted central server. This paper shows that a significantly weaker assumption suffices: each pair of participants shares a randomness seed unknown to others. In a setting where malicious participants may collude with an untrusted server, we propose CafCor, an algorithm that integrates robust gradient aggregation with correlated noise injection, using shared randomness between participants. We prove that CafCor achieves strong privacy-utility trade-offs, significantly outperforming local differential privacy (DP) methods, which do not make any trust assumption, while approaching central DP utility, where the server is fully trusted. Empirical results on standard benchmarks validate CafCor’s practicality, showing that privacy and robustness can coexist in distributed systems without sacrificing utility or trusting the server.
Differential Privacy, Robustness, Distributed Optimization, Federated Learning
@@ 1 Introduction
Differential privacy (DP) (Choudhury et al., 2019; Hu et al., 2020; Noble et al., 2022) and robustness (Yin et al., 2018; Karimireddy et al., 2021; Allouah et al., 2023b) have been extensively studied in isolation in distributed machine learning. However, their intersection, which is crucial for building trustworthy learning systems, remains underexplored. Recent work shows that combining local DP (LDP), which assumes no trust, with robustness incurs a fundamental utility cost (Allouah et al., 2023c). This reinforces earlier findings that ensuring privacy under LDP, even without malicious workers, is prohibitive in terms of utility (Duchi et al., 2013). In contrast, the central DP (CDP) model, where the server is fully trusted, enables significantly better utility than LDP, particularly as the number of workers increases (Liu et al., 2021; Hopkins et al., 2022). These findings highlight the trade-offs between trust, privacy, and robustness in distributed learning, underscoring the need for alternative approaches that mitigate these limitations.
@@ 1.2 Related Work
Most works combining DP and robustness focus on local privacy models, often only achieving per-step privacy guarantees and suboptimal learning performance (Guerraoui et al., 2021; Zhu & Ling, 2022; Xiang & Su, 2022; Ma et al., 2022). Others reveal a three-way trade-off between privacy, robustness, and utility for discrete distribution estimation in non-interactive settings (Cheu et al., 2021; Acharya et al., 2021) and more general interactive learning tasks (Allouah et al., 2023c). In the CDP model where the server is fully trusted, the privacy-utility trade-off is far superior to the LDP setting (Duchi et al., 2013, 2018). Prior work on centralized robust mean estimation with DP (Liu et al., 2021; Hopkins et al., 2022; Alabi et al., 2023) achieves near-optimal error bounds, disentangling the three-way trade-off by matching the optimal sample complexity in both private and robust settings. Empirical studies confirm this advantage in standard federated learning tasks (Xie et al., 2023).
Intermediate threat models between local and central DP, such as those leveraging cryptographic primitives or anonymization via shuffling (Jayaraman et al., 2018; Erlingsson et al., 2019; Cheu et al., 2019), can approach the privacy-utility trade-off of central DP but often incur high computational costs, rely on trusted entities, or lack robustness. Correlated noise schemes have also been explored for privacy without a trusted server; early methods like secret sharing (Shamir, 1979) ensure input security but do not privatize the output, leaving it vulnerable to inference attacks (Melis et al., 2019). More refined approaches (Imtiaz et al., 2021) combine correlated and uncorrelated Gaussian noise to protect averages, while decentralized extensions remain limited to averaging tasks (Sabater et al., 2022) and lack robustness (Allouah et al., 2024). We also note that (Sabater et al., 2022) consider malicious workers but do not provide utility guarantees in their presence, and only use cryptographic checks, e.g., to verify boundedness of all messages. On the other hand, they consider sparse topologies which induces lower communication costs, unlike our quadratic communication complexity for seed exchange, which we recall is negligible since it happens once before training and only includes integers, compared to the large-scale models exchanged iteratively during training. Recently, Gu et al. (2023) considered malicious workers and an untrusted server but relied on computationally intensive tools like zero-knowledge proofs and secure aggregation, whereas our method avoids these complexities. Moreover, they do not quantify the utility cost of their solution.
@@ 2.3 Assumptions
Our results are derived under standard assumptions. Data heterogeneity is modeled via Assumption 2.1 (Allouah et al., 2023c), which subsumes standard assumptions in federated learning (Karimireddy et al., 2020) by bounding gradient variance in all directions. To establish the convergence guarantees of CafCor, we adopt the standard Assumption 2.2 on the variance of stochastic gradients. Assumption 2.3 ensures bounded gradients at any data point, a common requirement in differentially private optimization to avoid complications due to clipping (Agarwal et al., 2018).
@@ 4.1 Privacy Analysis
The proof of Theorem 4.1, deferred to Appendix A, relies upon computing the Rényi divergence induced by the correlated noise scheme, before converting back to DP following standard conversion results (Mironov, 2017). Indeed, we derive the exact expression of the aforementioned Rényi divergence in Lemma A.4, which we use directly in our Rényi-based privacy accountant for SecLDP in practice. This exact computation was not feasible in previous related works, as they either considered different communication topologies (Sabater et al., 2022; Allouah et al., 2024). The major difference with the aforementioned analyses is granularity: we consider $0{\leq}q{\leq}f$ workers who may poison training but are not curious, while prior works considered one extreme or the other, i.e., $q\in\{0,f\}$. This granularity induces improved empirical performance (Figure 1).
@@ 4.2 Utility Analysis
Proposition 4.1 establishes an important property satisfied by CAF. The above corresponds to the high-dimensional robustness criterion studied by Allouah et al. (2023c), shown to be stronger than existing ones in the distributed learning literature, e.g., that of Karimireddy et al. (2022). Furthermore, we recall from Allouah et al. (2023c) that standard robust aggregations, such as the trimmed mean and median (Yin et al., 2018), can only achieve a weaker variant of Proposition 4.1, where the right-hand side of the inequality involves the trace instead of the maximum eigenvalue operator. Since the trace of a $d\times d$ matrix can be up to $d$ times its largest eigenvalue, standard robust aggregations only satisfy Proposition 4.1 with an additional multiplicative factor of $d$ in $\kappa$ (4.1), leading to a looser robustness bound.
Moreover, the so-called SMEA aggregation of Allouah et al. (2023c) also satisfies Proposition 4.1 up to constants. However, SMEA is computationally impractical as its time complexity grows exponentially with $f$. In contrast, CAF is efficient, running in polynomial time $\mathcal{O}(f(nd^{2}+d^{3}))$. Indeed, at each iteration, at least one input ($i\in\operatorname*{argmax}_{j\in[n]}\tau_{j}$) is assigned a weight of zero. Thus, CAF terminates within at most $2f$ iterations, each requiring the computation of the top eigenvector and the formation of the empirical covariance matrix.To further improve efficiency, we implement a CAF variant with reduced complexity $\mathcal{O}(fnd\log{d})$ by approximating the top eigenvector using the power method. This approach avoids explicit covariance matrix formation and instead relies on matrix-vector products. We adopt this more efficient variant in our experiments in Section 5, as it can be easily shown to satisfy Proposition 4.1 up to constants.
@@
Our result recovers the state-of-the-art convergence analysis in robust optimization in the local privacy model (Allouah et al., 2023c) when setting $\sigma_{\mathrm{cor}}=0$. The key difference with the latter analysis is the role of the correlated noise magnitude $\sigma_{\mathrm{cor}}$, which appears in the dominating term inside $\overline{\sigma}$, in both strongly convex and non-convex settings. Moreover, our convergence rates exhibit the standard dependence on the number of iterations in smooth stochastic optimization: $\mathcal{O}(\tfrac{1}{T})$ for the strongly convex case and $\mathcal{O}(\tfrac{1}{\sqrt{T}})$ for the non-convex case. The presence of a non-vanishing term in $T$ is fundamental in robust optimization in heterogeneous data settings (Karimireddy et al., 2022). Finally, our upper bound in the strongly convex case also applies to non-convex loss functions that satisfy the Polyak-Lojasiewicz inequality (Karimi et al., 2016).
@@ 6 Conclusion
First, we remark that Algorithm 2 terminates after at most $2f$ iterations. Indeed, at each iteration, at least one weight $c_{i},i\in[n]$, becomes zero—this is true for any $i\in[n]$ such that $i\in\operatorname*{argmax}_{j\in[n]}\left\langle\hat{v}_{c},\,x_{j}-\hat{\mu}_{c}\right\rangle^{2}$—for the remaining iterations. Therefore, since the weights $c_{i},i\in[n],$ are all in $[0,1]$, it is guaranteed that $\sum_{i=1}^{n}c_{i}<n-2f$ after $12f+$ 1 iterations.
@@ C.2 Proof Outline
Our analysis of CafCor (Algorithm 1), is inspired from standard analyses in non-private robust distributed learning (Karimireddy et al., 2021; Farhadkhani et al., 2022; Allouah et al., 2023c). We include all proofs for completeness.
@@
The dimension factor $d$ due to correlated and uncorrelated noises is divided by $n-f$, which would not have been possible without leveraging the Gaussian nature of the noise. To do this, we use a concentration argument on the empirical covariance matrix of Gaussian random variables, stated in Lemma C.6. Moreover, the correlated noise magnitude $\sigma_{\mathrm{cor}}^{2}$ has a multiplicative factor $n$, which essentially reflects the fact that a correlated noise term is added per worker, for each worker. Finally, the remaining term $G_{\mathrm{cov}}^{2}$ of the upper bound is only due to data heterogeneity, and also appears in Allouah et al. (2023c). Similar to the aforementioned work, an important distinction from Karimireddy et al. (2022) is that $G_{\mathrm{cov}}^{2}$ is a tighter bound on heterogeneity, compared to $G^{2}$ the bound on the average squared distance from 2.1. This is because the drift $\Delta_{t}$ is not an average squared distance, but rather a bound on average squared distances of every projection on the unit ball. Controlling this quantity requires a covering argument (stated in Lemma C.4).
@@
We prove Theorem 4.2 in the strongly convex case in Section C.3.1, and in the non-convex case in Section C.3.2. These proofs follow along the lines of Allouah et al. (2023c), and we include them here for completeness.
@@ C.6 Proof of Supporting Lemmas
Before proving Lemmas C.1, C.2 and C.3 in Sections C.6.2, C.6.3 and C.6.4, respectively, we first present some additional results in Section C.6.1 below. Most of the proofs follow along the lines of Allouah et al. (2023c), and we include them here for completeness.
@@ D.1 CafCor Under Three Privacy Threat Models
Comparison with SMEA (Allouah et al., 2023c)
We also included in our experiments a comparison with the SMEA robust aggregation method on the MNIST dataset (see Figures 7, 8, and 9). These results show that SMEA matches CAF’s accuracy, significantly outperforming other standard robust methods, and even approaching the averaging baseline. This observation aligns with our theoretical findings (see Proposition 4.1 and subsequent discussion).
However, we emphasize that SMEA incurs a substantial computational cost, which further supports the empirical advantage of CAF, despite the two methods offering similarly strong robustness in practice. To highlight SMEA’s prohibitive computational complexity, we report runtime ratios relative to simple averaging (lower is better), with $n=30$ workers and $f=3$ malicious workers, on models of varying dimension $d$.
Dimension ($d$) SMEA CAF Meamed GM
$0.25\times 10^{7}$ 30,251 28 112 62
$0.5\times 10^{7}$ 61,142 51 197 78
$1\times 10^{7}$ 117,255 100 378 126
Table 1: Computation time (relative to simple averaging) across different dimensions for various aggregation methods.
As seen in Table 1, CAF dramatically outperforms SMEA in terms of runtime, and also maintains a clear advantage over other robust aggregation baselines such as Meamed and GM. Combined with CAF’s superior performance on test accuracy, these results underscore CafCor’s practical relevance for trustworthy federated learning.